### v1

# Detecting Multimedia Generated by Large AI Models: A Survey

Abstract
The rapid advancement of Large AI Models (LAIMs), particularly diffusion models and large language models, has marked a new era where AI-generated multimedia is increasingly integrated into various aspects of daily life. Although beneficial in numerous fields, this content presents significant risks, including potential misuse, societal disruptions, and ethical concerns. Consequently, detecting multimedia generated by LAIMs has become crucial, with a marked rise in related research. Despite this, there remains a notable gap in systematic surveys that focus specifically on detecting LAIM-generated multimedia. Addressing this, we provide the first survey to comprehensively cover existing research on detecting multimedia (such as text, images, videos, audio, and multimodal content) created by LAIMs. Specifically, we introduce a novel taxonomy for detection methods, categorized by media modality, and aligned with two perspectives: pure detection (aiming to enhance detection performance) and beyond detection (adding attributes like generalizability, robustness, and interpretability to detectors). Additionally, we have presented a brief overview of generation mechanisms, public datasets, and online detection tools to provide a valuable resource for researchers and practitioners in this field. Furthermore, we identify current challenges in detection and propose directions for future research that address unexplored, ongoing, and emerging issues in detecting multimedia generated by LAIMs. Our aim for this survey is to fill an academic gap and contribute to global AI security efforts, helping to ensure the integrity of information in the digital realm. The project link is https://github.com/Purdue-M2/Detect-LAIM-generated-Multimedia-Survey.
Index Terms:
Media Forensics, Deepfake, Detection, Large AI Models, Diffusion Models, Large Language Models, Generation
1 Introduction
Large AI Models (LAIMs) are characterized by their exceptionally high parameter counts, often reaching billions, as highlighted in the work of [1, 2, 3]. These models typically include Diffusion Models (DMs) and Large Language Models (LLMs), both of which are trained on extensive datasets and require significant computational resources. A key distinction of LAIMs, especially DMs, from traditional AI models like Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs), lies in their scalability and generation quality. More specifically, LAIMs excel in tackling intricate tasks such as language comprehension, pattern recognition, and the generation of highly realistic multimedia content encompassing text, images, videos, and audio. Since 2020, as documented by [2], these models have attracted substantial interest due to their consistent improvements in performance and versatility across various domains, including natural language processing [1], computer vision [4], and health informatics [3].
Recent advances in LAIMs offer numerous benefits for humans in various aspects of life and work. Specifically, the benefits of LAIM-generated multimedia span a wide range of fields, from enhancing education [5] and healthcare [6] to boosting creativity, and optimizing business processes [7], and improving accessibility [8]. These advances improve task efficiency and unlock new avenues for innovation and problem-solving.
Meanwhile, multimedia generated by LAIMs also carries significant risks of misuse and societal upheaval, as discussed in [9, 10]. Specifically, these models can be employed to craft convincing fake news, deepfakes (highly realistic fake images, audios, and videos created using AI algorithms [11]), and other forms of misinformation, thereby challenging the information accuracy and public trust [12, 13, 14]. For instance, LAIM-generated multimedia can be weaponized for political propaganda or manipulative advertising, exploiting its ability to create persuasive and tailored content. There are also ethical dilemmas regarding the use of LAIMs to produce art or content that imitates human creativity, sparking debates over originality, intellectual property rights, and the intrinsic value of human artistic expression [15]. Furthermore, LAIMs’ ability to generate multimedia could potentially impact employment in creative fields, fueling concerns about the displacement of human workers in journalism, the arts, and entertainment. A notable example of this tension was the months-long strike in the summer of 2023, where writers and performers fought against major Hollywood studios [12].
Such challenges have highlighted the urgent need for effective detection methods for multimedia produced by LAIMs. In recent years, there has been a significant increase in research focused on this area. However, to the best of our knowledge, there is a notable lack of systematic surveys specifically addressing the detection of LAIM-generated multimedia, in contrast to the numerous surveys focusing on multimedia generation using LAIMs. To bridge this gap, we present this first comprehensive survey, which not only fills a critical academic void but also aligns with AI security initiatives of various governments worldwide, such as the AI Safety Summit 2023 [16] and the United States government’s “AI Executive Order” [17].
In this survey, we provide a detailed and thorough review of the existing research on identifying multimedia generated by LAIMs, particularly emphasizing Diffusion Models and Large Language Models. Our goal is to guide researchers toward understanding the current challenges and exploring potential future directions in this field, aiming to reinstate trust in digital content among users. Furthermore, we endeavor to show that, despite the high degree of realism in LAIM-generated multimedia, it can still be identified, which is crucial for its ethical use and for maintaining the integrity of information in the digital world. The dynamic and ongoing interplay between generating and detecting multimedia using LAIMs is shown in Fig. 1.
Figure 1: A cat-and-mouse game between generating and detecting multimedia ( text , image , video , audio , and multimodal ) using LAIMs, showcasing only representative works. Q1 represents from Jan to Mar, Q2: Apr-Jun, Q3: Jul-Sep, Q4: Oct-Dec.
1.1 Related Works
While there are a few surveys [18, 19, 20, 21, 22] addressing the detection of multimedia generated by LAIMs, like Diffusion Models (DMs) and Large Language Models (LLMs), their scopes are very limited. Specifically, surveys like [18, 19, 20, 21] mainly concentrate on the detection of LAIM-generated text, overlooking other multimedia forms such as images, videos, and audios. Additionally, while these surveys provide insights into detection techniques, they tend to focus solely on detection mechanisms without delving into broader aspects or applications of these technologies beyond detection, which are discussed in our survey.
The survey by Cardenuto et al. [22] falls short of being comprehensive. They present a high-level overview of challenges and opportunities in generation and detection using a broad range of AI models. However, their focus is not specifically on large AI models, leading to an oversight of numerous existing detection works. Additionally, their survey lacks in-depth discussion about datasets in the context of detection. Crucially, the detection methods discussed in [22] are confined to forensic types and do not delve into the specifics of detection models and methodologies, which is one of the central focuses of our survey. They also overlook detection approaches that utilize multimodal data modality, an area our survey encompasses.
In summary, there is currently no comprehensive survey covering the detection of fake media in text, images, videos, audio, and multimodal content, especially from the perspective of large AI models. Our survey attempts to fill this gap by making the following contributions.
1.2 Contributions
  1. 1.
    This is the first comprehensive survey on detecting multimedia generated by LAIMs, covering text, images, videos, audio, and multimodal content. Through this organized and systematic survey, we hope to help push research in this field forward.
  2. 2.
    We provide a brief overview of LAIMs, their generation mechanisms, and the content they generate, shedding light on the current status of the cat (generation)-and-mouse (detection) game in this field. Additionally, we present public datasets tailored for detection tasks.
  3. 3.
    This survey reviews the most up-to-date detection methodologies for LAIM-generated multimedia. We innovatively categorize them for each modality into two main groups: pure detection and beyond detection. This taxonomy offers a unique perspective that has not been previously explored. Within these two categories, we further classify the methods into more specific subcategories based on their common and distinct characteristics.
  4. 4.
    We have conducted thorough examinations of online detection tools. This provides a valuable resource for both researchers and practitioners in the field. Additionally, we pinpoint current challenges faced by recent detection methods and provide directions for future work.
The remained of this survey is organized as follows: In Section 2, we briefly introduce LAIMs, including generation mechanisms and datasets. In Section 3, we classify detection by their functionality and organize them into our newly defined taxonomy. We summarize online detection tools in Section 4. Critical challenges faced by detectors and potential future directions are discussed in Section 5. Finally, we conclude the survey in Section 6.
Figure 2: Illustrations of different types of multimedia generation process based on LAIMs.
Modality Dataset Content Link I2O #Real #Generated Source of Real Media Generation Method Year
Text Stu.Essays[23] Essays Link T2T 1,000 6,000 IvyPanda [24] ChatGPT 2023
Writing[23] Essays Link T2T 1,000 6,000 Reddit WritingPrompts [25] ChatGPT 2023
News[23] Essays Link T2T 1,000 6,000 Reuters 50-50 [26] ChatGPT 2023
Paraphrase[27] Essays Link T2T 98,280 163,710 Arxiv, Wikipedia, Theses GPT-3, T5[28] 2022
AA[29] Essays Link T2T 1,064 8,512 News Media GPT-1&2, CTRL[30], GROVER[31] 2020
OUTFOX [32] Essays Link T2T 15,400 15,400 Feedback Prize [33] ChatGPT, GPT-3.5, T5 [28] 2023
MULTITuDE[34] News Link T2T 7,992 66,089 MassiveSumm [35] GPT-3&4, ChatGPT 2023
TuringBench[36] News Link T2T 8,854 159,758 News Media GPT-1&2&3, CTRL [30], GROVER [31] 2021
GPTUnmixed[37] News Link T2T 5,454 5,454 News Media GPT-3.5 2023
GPTMixed[37] News Link T2T 5,032 5,032 News Media GPT-3.5 2023
GPABench[38] Writing Link T2T 150,000 450,000 Arxiv GPT-3.5 2023
HPPT[39] Abstracts Link T2T 6,050 6,050 ACL Anthology [40] ChatGPT 2023
TweepFake [41] Tweets Link T2T 12,786 12,786 GitHub, Twitter GPT-2, RNN, LSTM 2021
SynSciPass [42] Passages Link T2T 99,989 10,485 Scientific papers GPT-2, BLOOM[43] 2022
DFText[44] General Link T2T 154,078 294,381 Reddit, ELI5[45],Yelp, XSum [46] GPT, GLM [47], LLaMA, T5 [28] 2022
HC-Var[48] General Link T2T 90,096 45,000 XSum [46], IMDb, Yelp, FiQA [49] ChatGPT 2023
HC3 [50] General Link T2T 26,903 58,546 FiQA[49], ELI5[45], Meddialog[51] ChatGPT 2023
M4[52] General Link T2T 32,798 89,683 WikiHow [53], Arxiv, Reddit ChatGPT, LLaMA,T5 [28], BLOOM [43] 2023
MixSet[54] General Link T2T 300 3,600 Email[55], BBC News[56], ArXiv
GTP-4, LLaMA2 [57]
2024
InternVid[58] Captions Link V2T 7,000,000 234,000,000 YouTube ViCLIP [58] 2023
Image DFF[59] Face Link T/I2I 30,000 90,000 IMDB-WIKI[60] SDMs, InsightFace [61] 2023
RealFaces [62] Face Link T2I 258 25,800 Prompts SDMs 2023
OHImg [63] Overhead Link T/I2I 6,475 6,675 MapBox [64], Google Maps GLIDE [65], DDPM [66] 2023
Western Blot[67] Biology Link I2I $\sim$14,000 $\sim$24,000 Western Blot DDPM, Pix2pix [68], CycleGAN [69] 2022
Synthbuster[70] General Link T2I - 9,000 Raise-1k[71] DALL·E 2&3, Midjourney, SDMs, GLIDE[65] 2023
GenImage[72] General Link T/I2I 1,331,167 1,350,000 ImageNet SDMs, Midjourney, BigGAN [73], 2023
CIFAKE[74] General Link T2I 60,000 60,000 CIFAR-10 SD-V1.4 2023
AutoSplice[75] General Link T2I 2,273 3,621 Visual News [76] DALL·E 2 2023
DiffusionDB[77] General Link T2I 3,300,000 16,000,000 DiscordChatExporter [78] SD 2023
ArtiFact[79] General Link T/I2I 964,989 1,531,749 COCO, FFHQ [80], LSUN SDMs, DDPM [66],LDM [81], CIPS [82] 2023
HiFi-IFDL[83] General Link T/I2I $\sim$600,000 1,300,000 FFHQ [80], COCO, LSUN DDPM [66], GLIDE [65], LDM [81],GANs 2023
DiffForensics[84] General Link T/I2I 232,000 232,000 LSUN, ImageNet LDM [81], DDPM [66], VQDM [85], ADM [86] 2023
CocoGlide[87] General Link T2I 512 512 COCO GLIDE[65] 2023
LSUNDB[88] General Link T/I2I 250,000 250,000 LSUN DDPM [66], LDM [81], StyleGAN [80] 2023
UniFake[89] General Link T2I 8,000 8,000 LAION-400M [90] LDM [81], GLIDE [65] 2023
REGM [91] General Link T/I2I - 116,000 CelebA [92], LSUN 116 publicly available GMs 2023
DMimage [93] General Link T2I 200,000 200,000 COCO, LSUN LDM [81] 2022
AIGCD[94] General Link T/I2I 360,000 508,500 LSUN, COCO, FFHQ [80] SDMs, GANs,ADM [86],DALL·E 2,GLIDE[65] 2023
DIF [95] General Link T/I2I 84,300 84,300 LAION-5B [96] SDMs, DALL·E 2, GLIDE [65], GANs 2023
Fake2M[97] General Link T/I2I - 2,300,000 CC3M[98] SD-V1.5 [99], IF [100], StyleGAN3 2023
Video DiffHead[101] Face Link I.A2V - 820 CREMA [102] Diffused Heads: build on DDPM 2023
Audio LibriSeVoc[103] Speech Link T2A 13,201 79,206 LibriTTS[104] DiffWave [105], WaveNet [106] 2023
Multi- modal $DGM^{4}$ [14] News Link T/I2T 77,426 152,574 Visual News [76] B-GST [107], StyleCLIP [108], HFGI [109] 2023
COCOFake[110] General Link T/I2T 113,287 566,435 COCO SDMs 2023
TABLE I: Summary of public datasets that are generated by LAIMs. I2O: Input-to-Output, T2T: Text-to-Text, V2T: Video-to-Text, T/I2I: Text-to-Image,Image-to-Image, T2A: Text-to-Audio, I.A2V: (Image conditioned with Audio)-to-Video. Only representative works are listed in ”Source of Real Media” and “Generative Method”.
2 Generation
In this section, we provide an overview of large generative AI models, their generation mechanisms, and the type of content they generate.
$\clubsuit$ Text. Machine-generated text, primarily driven by the advent of LLMs, is increasingly permeating many aspects of our daily lives. The exceptional proficiency of LLMs in understanding, following, and complex reasoning [111] has established their dominance in text creation. Recently, we have witnessed a surge of LLMs including OpenAI GPT-4 [112], Google Gemini [113], Meta LLaMA2 [57] as well as the remarkable performance they have achieved in various tasks, such as News [31, 36], Question Answering [45], Biomedicine [114], Code Generation [115], Tweets [41], and Scientific Writing [42], see Fig. 2 a). More details can be found in [116, 1].
Datasets. The prevalent datasets for LLM-generated text detection are listed in Table I. For example, HC3 [50] stands as one of the initial open-source efforts to compare ChatGPT-generated text with human-written text. Due to its pioneering contributions in this field, the HC3 corpus has been utilized as a valuable resource in numerous subsequent studies. Moreover, Authorship Attribution [29] and TuringBench[36] datasets are proposed to evaluate the attribution capability of detectors. HPPT [39] can benchmark task for detecting ChatGPT-polished texts.
$\clubsuit$ Image. In the challenging task of image synthesis, diffusion models (DMs) [66, 117, 118] have emerged as the new state-of-the-art family of deep generative models. The image generation process in DMs usually contains two processes [66]: a forward process that progressively destroys data by adding noise and a reverse process that learns to generate new data by denoising. More details can be found in [4, 119]. Current research on diffusion models is mostly based on three predominant formulations: denoising diffusion probabilistic models (DDPMs) [66], score-based generative models (SGMs) [117], and stochastic differential equations (Score SDEs) [118]. Building upon them, more advanced models have emerged in image generation, including OpenAI DALL·E 3 [120], Stable Diffusion V2 [81], Google Imagen2 [121], Midjourney [122], Amazon Titan Image Generator [123], and Meta Emu Edit [124], see Fig. 2 b).
Datasets. There are several million-scale generated images detection datasets: GenImage [72], DiffusionDB [77], ArtiFact [79], HiFi-IFDL [83], and Fake2M [97]. DiffusionDB stands out as the first large-scale Image-Prompt pairs dataset, though its image is only generated by Stable Diffusion. GenImage, ArtiFact, HiFi-IFDL, and Fake2M contain images generated by various DMs and GANs. Unique to HiFi-IFDL is its inclusion of high-resolution forgery masks, making it particularly useful for both detection and localization tasks. Refer Table I for more details.
$\clubsuit$ Video. In the pursuit of high-quality video generation, recent research has also turned to diffusion models. Early work [125, 126] are based on DDPM [66] for video generation. Current research extends text-to-image (T2I) diffusion models to text-to-video (T2V) generation. Meta proposes Make-A-Video [127] extending a diffusion-based T2I model to T2V through spatiotemporally factorized diffusion model. A more advanced model, Emu Video [128], is their latest video generation milestone. Unlike Make-A-Video, which requires five models, Emu Video is a simple unified architecture for video generation and uses just two diffusion models. Given a text prompt, Google’s Imagen Video [129] generates high-resolution videos using a base video generation model and a sequence of interleaved spatial and temporal video super-resolution models. More recently, Stability AI’s Stable Video Diffusion [130] and Runway’s GEN-2 [131] offer latent video diffusion models for video generation, while Pika [132] introduces a platform aimed at broadening creative possibilities in video creation, as shown in Fig. 2 c).
Datasets. Though building upon the success of T2I generation, T2V generation requires temporally smooth and realistic motion that matches the text, besides high-quality visual content, making it still in the nascent stage compared with image generation. Furthermore, video generation is much more computationally costly than image synthesis. To our best knowledge, only one work [101] contributes their generated talking head videos via LAMs, see Table I.
$\clubsuit$ Audio. Most audio synthesis by diffusion models focuses on text-to-speech (TTS) tasks. One line of work [133, 134, 105] first generates acoustic features, e.g., mel-spectrogram, and then outputs waveform with a vocoder. Another branch of work [135, 136] attempts to solve the TTS task in an end-to-end manner, as shown in Fig. 2 d). Diff-TTS [133] is the first work that applies DDPM [66] to mel-spectrogram generation. It transforms the noise signal into a mel-spectrogram via diffusion time steps. [134, 105] apply diffusing models to vocoder for generating waveform based on mel-spectrogram. Instead of treating acoustic modeling and vocoder modeling as independent processes,  [135, 136] generate audio from text without acoustic features as explicit representation.
Datasets. Due to comparatively less research attention on the detection of fake audio, LibriSeVoc [103] in Table I is the only public dataset that includes diffusion model-based vocoders (WaveGrad [134] and DiffWave [105]) so far.
$\clubsuit$ Multimodal. Multimodal learning refers to an embodied learning situation that includes learning multiple modalities such as text, image, video, and audio [137]. From a generation perspective, visual generation tasks, such as text-to-image and text-to-video, are regarded as multimodal generation. The generative models for these tasks are trained to learn visual representations and language understanding for visual generation. From the detection aspect, detectors that learn multiple modalities for forgery detection are categorized into multimodal. In this context, we define "multimodal generation" from a detection perspective, referring to frameworks capable of creating multimodal output. Multimodal generation process normally contains foundation models as encoders (e.g., CLIP [138], ViT [139] ) and decoders (e.g., Stable Diffusion [81] ), and a LLM for taking language-like representations from encoders for semantic understanding and produces modality signal tokens for guiding content output, see Fig. 2 e). Most recent work, such as HuggingGPT [140], AudioGPT [141], and NExT-GPT [142], all are based on the language understanding and complex reasoning capabilities of LLMs, and utilizing existing off-the-shelf multimodal encoders and decorders as tools to execute multimodal input and output.
Datasets. The multimodal datasets in Table I contain multi-types of fake media. For example, $DGM^{4}$ [14] and COCOFake [110] includes synthesized text and image.
3 Detection
{forest}
[…]
3), basic, l sep=5mm, [Multi-
modal (§3.5), xnode, l sep=5mm, [Beyond
3.5.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=31.2mm, [Coccomini et al. [143],VERITE [144]   , 5mmnode]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ Frequency-based   , 4mmnode,l sep=5mm, [UFAFormer [145]   , 5mmnode]] [ Spatial-based   , 4mmnode,l sep=5mm, [HAMMER [14], HAMMER++ [14], Wang et al. [146]   , 5mmnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=31.2mm, [Xu et al. [147]   , 5mmnode]] [$\clubsuit$ Generalization  , nnode,l sep=5mm, [ Contrastive Learning   , 4mmnode,l sep=5mm, [LASTED [148]   , 5mmnode]] [ Prompt Tuning   , 4mmnode,l sep=5mm, [AntifakePrompt [149]   , 5mmnode]] ] [$\clubsuit$ Attribution  , nnode,l sep=31.2mm, [De-Fake [150]   , 5mmnode]]] [Pure
3.5.1)  , tnode,l sep=5mm, [$\clubsuit$ Text-image Inconsistency  , nnode,l sep=31.2mm, [DIDAN [13]   , 5mmnode]] [$\clubsuit$ Text-assisted  , nnode,l sep=31.2mm, [Amoroso et al. [151]   , 5mmnode]] ] ] [Audio
3.4), xnode, l sep=5mm, [Pure
3.4.1)  , tnode, l sep=5mm, [$\clubsuit$ Vocoder-based  , nnode,l sep=31.2mm, [DetectVocoder [103]   , 5anode]]] ] [Video
3.3), xnode, l sep=5mm, [Beyond
3.3.1)  , tnode, l sep=5mm, [$\clubsuit$ Generalization  , nnode,l sep=31.2mm, [RevisitVideo [152]   , 5vnode]]] ] [Image
3.2), xnode, l sep=5mm, [Beyond
3.2.2)  , tnode,l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=31.2mm, [Corvi et al. [93, 153], Ricker et al. [88], Cocchi et al. [110], Papa et al. [62], Porcile et al. [154], Mandelli et al. [67], Carriere et al. [155]   , 5inode]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ Frequency-based   , 4inode,l sep=5mm, [D4 [156]   , 5inode]] [ Spatial-based   , 4inode,l sep=5mm, [GLFF [157], Xu et al. [158], Local Statistics [159]   , 5inode]]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ Weakly-supervised   , 4inode,l sep=5mm, [Tantaru et al. [160]   , 5inode]] [ Fully-supervised   , 4inode,l sep=5mm, [IFDL [83], PAL [161], TruFor [87]   , 5inode]]] [$\clubsuit$ Interpretability  , nnode,l sep=31.2mm, [Aghasanli et al. [162]   , 5inode]] [$\clubsuit$ Generalization  , nnode,l sep=31.2mm, [Online Detection [163], Ojha et al. [89], Cozzolino et al. [164], LSDA [165], Fingerprintnet [166], FACTOR [167], Dogoulis et al. [168], DNF [169]   , 5inode]] [$\clubsuit$ Attribution and Model Parsing  , nnode,l sep=5mm, [ Model Parsing   , 4inode,l sep=5mm, [Asnani et al. [91]   , 5inode]] [ Attribution   , 4inode,l sep=5mm, [Guarnera et al. [170]   , 5inode]] ]] [Pure
3.2.1)  , tnode, l sep=5mm, [$\clubsuit$ Frequency-based Methods  , nnode,l sep=31.2mm, [Wavelet-Packets [171], AUSOME [172], Xi et al. [173], Synthbuster [70]   , 5inode]] [$\clubsuit$ Spatial-based Methods  , nnode,l sep=31.2mm, [Zhong et al. [94], Nguyen et al. [174], multiLID [175]   , 5inode]] [$\clubsuit$ Diffuser Fingerprints   , nnode,l sep=31.2mm, [DIF [95], DIRE [84], SeDID [176]   , 5inode]] [$\clubsuit$ Physical/Physiological  , nnode,l sep=31.2mm, [Borji [177], Perspective [178], Lighting [179]   , 5inode]] ] ] [Text
3.1), xnode, l sep=5mm, [Beyond
3.1.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=5mm, [ Paraphrase Detection  , 4tnode,l sep=5mm, [Wahle et al. [27], Becker et al. [180]   , 5tnode]] [ Attribution  , 4tnode,l sep=5mm, [Antoun et al. [181]   , 5tnode]] [ Human Evaluation  , 4tnode,l sep=5mm, [ Guo et al. [50]   , 5tnode]] [ Generalization/Robustness  , 4tnode,l sep=5mm, [ChatLog [182], Zero-Shot [183], Training-based [48], Sarvazyan et al. [184], In The Wild [44]   , 5tnode]]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ Stylistic/Consistency  , 4tnode,l sep=5mm, [J-Guard [185], PHD  [186]   , 5tnode]] [ Adversarial Learning  , 4tnode,l sep=5mm, [RADAR [187], OUTFOX [32]   , 5tnode]] [ Adversarial Data Augmentation  , 4tnode,l sep=5mm, [Yang et al. [39], Shi et al. [188], MGTBench [189]   , 5tnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=5mm, [ Polish Ratio  , 4tnode,l sep=5mm, [Yang et al. [39]   , 5tnode]] [ Shapley Additive Explanations  , 4tnode,l sep=5mm, [Mitrovic et al. [190], Check-GPT [38]   , 5tnode]] [ P-values  , 4tnode,l sep=5mm, [Kirchenbauer et al. [191]   , 5tnode]] [ N-gram Overlaps  , 4tnode,l sep=5mm, [DNA-GPT [192]   , 5tnode]]] [$\clubsuit$ Generalization  , nnode,l sep=5mm, [ Contrastive Learning  , 4tnode,l sep=5mm, [Conda [193]   , 5tnode]] [ Structured Search  , 4tnode,l sep=5mm, [Ghostbuster [23]   , 5tnode]]] [$\clubsuit$ Attribution  , nnode,l sep=5mm, [ Origin Tracing  , 4tnode,l sep=5mm, [Sniffer [194]   , 5tnode]] [ Style Representation  , 4tnode,l sep=5mm, [Few-Shot [195]   , 5tnode]] [ Perplexity  , 4tnode,l sep=5mm, [LLMDet [196]   , 5tnode]] [ Stylometric/Statistical  , 4tnode,l sep=5mm, [Linguistic Model [29], GPT-who [197]   , 5tnode]] [ Deep-learning Based  , 4tnode,l sep=5mm, [TuringBench [198], Contra-X [199], XLNet-FT [200], TopRoBERTa [201]   , 5tnode]]]] [Pure
3.1.1)  , tnode, l sep=5mm, [$\clubsuit$ Hard-explainable Methods  , nnode,l sep=5mm, [ Positive-Unlabeled  , 4tnode,l sep=5mm, [ MPU [202]   , 5tnode]] [ Efficient Perturbations  , 4tnode,l sep=5mm, [ Efficient-DetectGPT [203], Fast-DetectGPT [204], DetectLLM [205]   , 5tnode]] [ Log Probabilities Curvature  , 4tnode,l sep=5mm, [DetectGPT [206]   , 5tnode]] [ Perplexity  , 4tnode,l sep=5mm, [HowkGPT [207], GPTZero [208]   , 5tnode]] ] [$\clubsuit$ Easy-explainable Methods  , nnode,l sep=5mm, [ Stylometry/Coherence  , 4tnode,l sep=5mm, [Stylometric Signals [209], Coherence Consistency [210]   , 5tnode]] [ Artifacts  , 4tnode,l sep=5mm, [Pu et al. [211]   , 5tnode]] [ Watermarking  , 4tnode,l sep=5mm, [Distillation-Resistant [212], Fernandez et al. [213], Robust Multi-bit [214], Christ et al. [215], Kuditipudi et al. [216], Unigram Watermark [217], Private Watermark [218]   , 5tnode]] ]] ] ]
Figure 3: Taxonomy of the literature on detecting multimedia generated by Large AI Models (LAIMs).
In this section, we provide a novel taxonomy of detection methods on LAIM-generated multimedia. An overview of the structure for this section is provided in Fig. 3. Specifically, we identify the functionality of these detectors and organize them into two categories for each data modality:
  1. 1.
    Pure Detection. The detection methods in this category only aim to improve detection performance.
  2. 2.
    Beyond Detection. It endows detectors extra characteristics (e.g., generalizability, robustness, interpretability) while keeping accurate and effective detection ability.
3.1 Text
There are existing surveys providing a detailed overview of current detection strategies and benchmarks [19, 18, 20]. Specifically, [18] categorizes the detection methods into three groups: training-based, zero-shot, and watermarking. Works in each group are further distinguished according to their detection scenarios such as black-box and white-box. The authors in [19] categorize the detection of LLM-generated content based on the techniques applied by detectors (e.g., adversarial learning, watermarking methods, and human-assisted methods) and the dependency of detectors on training (e.g., zero-shot detectors, fine-tuned detectors). [20] divides the existing literature into two parts: methods designed to detect LLM-generated text (e.g., watermarking-based, fine-tune-based, and zero-shot) and methods designed to evade detection (e.g., paraphrasing attacks, spoofing attacks). Although these organizational strategies apply to recent detection methods, their taxonomy may not be sufficient for adapting to new or evolving detection techniques. To this end, we provide a novel taxonomy based on Pure Detection and Beyond Detection.
Figure 4: Illustrations of pure detection methodologies for LAIM-generated text.
3.1.1 Pure Detection
We categorize the pure detection methodologies based on their comprehensibility toward the general populace, which can be identified as Easy-explainable and Hard-explainable methods, as shown in Fig. 4.
$\clubsuit$ Easy-explainable Methods. These approaches ensure that humans can straightforwardly understand the principles behind the detection technology. Such methods prioritize clarity and accessibility, making the technology approachable for non-specialist users.
Watermarking. Detecting watermarks in the text would be the most straightforward way for humans to understand the detection technique. Text watermarking injects algorithmically detectable patterns into the generated text while ideally preserving the quality and diversity of language model outputs. Most detection methods based on watermark require the watermark key for detection [212, 213, 214, 215, 216, 217], which are susceptible to security breaches and counterfeiting. [218] proposes the first method, which does not require the key during detection to alleviate this issue. They use the shared token embedding between the watermark generation and detection networks to improve the efficiency of training the detector.
Artifacts. While artifacts in text may not be as visually obvious as in images, this detection approach is simple and comprehensible to humans. Yet, intriguing artifacts or characteristics of LLM-generated texts can be elusive. [211] reveals those characteristics by perturbing these generated texts and offers the following empirical findings:
  • Artifacts exist mainly in the form of token co-occurrence and at the head of the vocabulary;
  • Content words contain more artifacts than stopwords;
  • High-level semantic/syntactic features contain much fewer artifacts than shallow features;
  • Some artifacts are present in higher-order N-grams, appearing as elements with unclear or vague meanings.
Stylometry/Coherence. Other works directly identify the stylistic changes and coherence inconsistency between LLM-generated texts and human-written ones. [209] employs stylometric analysis to distinguish between human and LLM-generated tweets by quantifying stylistic differences. [210] examines the linguistic structure of texts and uses the difference in coherence traced by entity consistency exhibited in LLM-generated text and human-written ones.
$\clubsuit$ Hard-explainable Methods. These methods involve detection techniques and analytical processes that may not be as easily comprehensible to most people but are quite accessible to researchers in the field.
Perplexity. This is a statistical metric used in language models. Several works employ perplexity to distinguish LLM-generated text from human-written text. For example, [207] discerns LLM-generated text, specifically homework assignments, by computing perplexity scores for student-authored and ChatGPT-generated responses. These scores then assist in establishing a threshold for discerning the origin of a submitted assignment. Moreover, the widely recognized tool, GPTZero [208], examines the text’s perplexity and burstiness metrics to estimate the likelihood of a review text being generated by LLMs for detection.
Log Probabilities Curvature. This refers to the logarithmic transformation of probabilities that a model assigns to sequences of words or tokens, and be used in certain research to detect LLM-generated text. Some works statistically utilize log probabilities to detect LLM-generated text, which require a certain level of knowledge among the general public regarding the generative mechanisms of LLMs to understand their detection methods. For example, the pioneer work DetectGPT [206] observes that LLM-generated text exhibits a more negative log probability curvature and leverages the curvature-based criterion based on random perturbations of the text, yielding promising results.
Efficient Perturbations. Recognizing the use of pure random perturbations in DetectGPT requires intensive computational cost, a series of works adopt various techniques to improve computational efficiency and enhance detection accuracy simultaneously. Specifically, [203] achieves similar performance with up to two times fewer queries than DetectGPT with a Bayesian surrogate model by selecting typical samples based on Bayesian uncertainty and interpolating scores from typical samples to other ones, making the perturbation process more focused and less resource-intensive. Bao et al. [204] increase the detection speed by 340 times by substituting DetectGPT’s perturbation step with a more efficient sampling step via conditional probability curvature. DetectLLM [205], another recent contribution, employs normalized perturbed log-rank for text detection generated by LLMs, asserting a lower susceptibility to the perturbation model and the number of perturbations compared to DetectGPT.
Positive-Unlabeled. Tian et al. [202] observe that LLM-generated short texts are highly similar to human, it is not suitable to assign these simple generated texts with either fake or real labels; rather, they are in an “Unlabeled” state. To this end, they model the task of LLM-generated text detection as a partial Positive-Unlabeled problem and formulate the Multiscale Positive-Unlabeled training framework to address the challenging task of short text detection without sacrificing long texts.
Figure 5: Illustrations of beyond detection methodologies for LAIM-generated text.
3.1.2 Beyond Detection
Detection Methods go beyond distinguishing between human and machine-generated content can be organized (see Fig. 5) as follows:
$\clubsuit$ Attribution. Determining which specific model may generate the test content, see Fig. 5 a).
Deep-learning Based. Most attribution models are deep-learning based [198, 199, 200, 201]. [200] addresses attribution by noting that synthetic texts carry subtle distinguishing marks inherited from their source models and these marks can be leveraged for attribution. TopRoBERTa [201] improves existing attribution solutions by capturing more linguistic patterns in LLM-generated texts by including a Topological Data Analysis (TDA) layer in the RoBERTa model. In this approach, RoBERTa RoBERTa captures contextual semantic and syntactic features, while a TDA layer analyzes the shape and structure of linguistic data.
Stylometric/Statistical. Besides the above deep learning-based classifier, some works solve the attribution task in a stylometric or statistical way. Stylometry is the statistical analysis of writing styles in texts. Uchendu et al. [29] propose a model named the Linguistic model. They train a Random Forest classifier with their proposed Authorship Attribution (AA) dataset and extract an ensemble of stylometric features, such as entropy and readability score. GPT-who [197] is a psycholinguistically-aware, statistical-based detector. It utilizes Uniform Information Density (UID) based features to model the unique statistical signature of each LLM and human author for accurate authorship attribution.
Perplexity. Inspired by the observation that perplexity serves as a reliable signal for distinguishing the source of generated text, [196] calculates the proxy perplexity to identify the sources from which the text is generated, such as Human, LLaMA, OPT [219], or others.
Style Representation. In pursuit of an approach that does not rely on samples from language models of concern at training time. Soto et al. [195] propose a few-shot strategy to detect LLM-generated text by leveraging representations of writing style estimated from human-written text.
Origin Tracing. This differs from the above model-wise attribution. It refers to tracing back a larger original model on which a smaller generative model is based. For example, trace text generated by Alpaca [220] back to ChatGPT and LLaMA. [194] introduces the first origin tracing tool, Sniffer, which utilizes the discrepancies between LLMs as features and then trains a simple linear classifier to help trace the origins.
$\clubsuit$ Generalization. Developing detectors with generalizability that can detect texts generated by generators never seen before, as shown in Fig. 5 b).
Structured Search. This method involves passing documents through weaker language models and then conducting a systematic search over possible combinations of their features. Verma et al. present Ghostbuster [23], a method for detection with generalizability based on structured search and linear classification. Ghostbuster runs the structured search to get the possible text features and then trains a classifier on the selected features to predict whether documents are AI-generated.
Contrastive Learning. [193] develops a contrastive domain adaption framework that blends standard domain adaption techniques with the representation power of contrastive learning to learn domain invariant representations for better generalizability.
$\clubsuit$ Interpretability. Exploring interpretable detectors that can provide explanations for detection results, see Fig. 5 c).
N-gram Overlaps. DNA-GPT [192] identifies GPT-generated text by exploiting the distinct text continuation patterns between human and AI-generated content. It provides evidence based on nontrivial N-gram overlaps to support explainable detection results.
P-values. [191] proposes a statistical test for detecting watermarks in text with interpretable P-values. P-values provide a statistically interpretable way to quantify the certainty that the detected pattern (i.e., watermark) in the text is not due to random chance.
Shapley Additive Explanations. Some researchers work toward gaining insight into the reasoning behind the model. For example, [190] fine-tunes a transformer-based model and uses it to make predictions, which are then explained using Shapley Additive Explanations (SHAP) [221], an eXplainable Artificial Intelligence (XAI) framework to extract explanations of the model’s decisions aiming at uncovering the model’s reasoning. Similarly, [38] also employs Shapley Values to compare the interpretations derived from word-level and sentence-level results.
Polish Ratio. Though the above studies provide interpretability, they do not address text that has been refined on a more granular level, such as ChatGPT-polished texts. To bridge this gap, [39] introduces a novel dataset termed HPPT (ChatGPT-polished academic abstracts) and proposes the “Polish Ratio" method, a measure of ChatGPT’s involvement in text based on editing distance. It calculates a similarity score between the original human-written texts and their polished versions, providing a more reasonable explanation for the detection outcome.
$\clubsuit$ Robustness. Developing detectors that can handle different attacks, see Fig. 5 d). In light of the vulnerability of detectors to different attacks and robustness issues, a significant body of research has been dedicated to utilizing adversarial learning as a mitigation strategy.
Adversarial Data Augmentation. [39, 188, 189] conduct the adversarial data augmentation process on LLM-generated text, their findings indicate that models trained on meticulously augmented data exhibited commendable robustness against potential attacks. The key technique in these methods involves employing adversarial attacks by creating misleading inputs, thereby enhancing the model’s competency to handle a wider range of situations where deception might be a factor.
Adversarial Learning. Besides these pre-processing methods, some works cast the spotlight on adversarial learning-based frameworks. Hu et al. [187] introduce RADAR, a framework for training robust detectors by simultaneously interacting with a paraphrasing model, which generates content aimed at evading detection. While effective against paraphrase attacks, its defense against other attacks is unexplored. Diverging from RADAR, Koike et al. [32] propose OUTFOX, which enhances robustness by interlinking detector and attacker outputs. The attacker learns from the detector’s predictions to create hard-to-detect essays, while the detector improves by learning from these adversarial essays. OUTFOX is particularly effective against paraphrase-based attacks.
Stylistic/Consistency. [185] studies and quantifies stylistic cues from the latent journalism process in real-world news organizations towards discriminating AI-generated news. Their J-Guard framework steers existing supervised AI text detectors while boosting adversarial robustness. The research by Tulchinskii et al. [186] explores consistent properties in human-written texts across different domains and skill levels. They estimate the geometry of text sample as an individual object and discover that real texts have a higher intrinsic dimension than artificial ones. This insight is used to estimate intrinsic dimensionality for detecting LLM-generated texts.
$\clubsuit$ Empirical Study. The empirical studies are crucial for advancing our understanding and capabilities in detecting LLM-generated texts, as shown in Fig. 5 e).
Generalization/Robustness. Tu et al. [182] observe a decline in RoBERTa-based detectors’ effectiveness over time through a month-long study of ChatGPT responses to long-form questions. Complementing this, Pu et al. [183] find that detectors trained on one generator can zero-shot generalize to another, especially when trained on a medium-size LLM for detecting content from a larger version.
A more comprehensive investigation on detectors’ generalization is conducted by Xu et al. [48]. They find that the trained models tend to overfit to some “irrelevant features” that are not principal for ChatGPT detection. This overfitting issue can be originated from the “incautious and insufficient” data collection process. Besides this, they provide an optimistic insight: the trained models are also capable of extracting “transferable features”, which are shared features that can help detect the ChatGPT generated texts from various topics and language tasks.
Additionally, [184] specifically investigate generalization of in-domain fine-tune detectors. They use the multi-domain, multilingual AuTexTification corpus to fine-tune various supervised detectors and discover that in-domain fine-tune detectors struggle against data from different models. Moreover, detection in the wild [44] conducts a more comprehensive experiment by building a wild testbed, involving 10 datasets covering diverse writing tasks and sources, and using 27 LLMs for creating texts. A key finding is the challenge of Out-Of-Distribution data in real-world scenarios, where the performance of detectors significantly declines, often barely better than random classification.
Human Evaluation. [50] conducts thorough human evaluations and linguistic analyses to compare the content generated by ChatGPT with that produced by humans. Key findings include: ChatGPT’s responses are more objective and formal, less likely to contain biased or harmful information, and generally longer and more detailed.
Attribution. An extensive investigation on model attribution, encompassing source model identification, model family classification, and model size classification, is conducted by Antoun et al. [181]. Their source model identification task involves classifying the text into one of the 50 LLMs, spanning various families and sizes. The results reveal several key findings: a clear inverse relationship between classifier effectiveness and model size, with larger LLMs being more challenging to detect, especially when the classifier is trained on data from smaller models.
Paraphrase Detection. Beyond detecting texts purely generated by LLMs and written by humans. Another line of empirical studies [27, 180] conduct extensive experiments on paraphrase detection concerning the severe threat to academic integrity. Specifically, [27] evaluates the performance of various detectors and perform a human study with 105 participants regarding their detection performance and the quality of generated examples. Their results suggest that humans struggle to identify large models-paraphrased text (53% mean accuracy). Human experts rate the quality of paraphrases generated by GPT-3 to be as high as that of original texts (clarity 4.0/5, fluency 4.2/5, coherence 3.8/5).  [180] evaluates various detection methods on several common paraphrase datasets, finding that human-authored paraphrases exceed machine-generated ones in terms of difficulty, diversity, and similarity.
3.2 Image
While DMs are evolving rapidly and producing increasingly realistic images, they still often make certain mistakes and leave identifiable fingerprints.
Figure 6: Illustrations of pure detection methodologies for LAIM-generated image.
3.2.1 Pure Detection
Research summarized here aims to identify DM-generated images by examining physical and physiological cues, as well as by focusing on enhancing detection accuracy.
$\clubsuit$ Physical/Physiological-based Methods. Physical-based methods detect DM-generated images by examining inconsistencies with real-world physics, such as lighting and reflections. Physiologically-based methods, on the other hand, investigate the semantic aspects of human faces [222], including cues such as symmetry, iris color, pupil shapes, skin, etc., see Fig. 6 a). These methods have stronger interpretability than data-driven methods, which have been widely adapted to detect GAN-generated images [223, 224].
Borji [177] outlines key cues for detecting DM-generated images that violate physical rules, as shown in Fig. 6 a) top. These cues include: i) Reflection. Generated images can exhibit artificial reflections that are inconsistent with the natural lighting and environment, such as those in glasses, mirrors, or pupils. ii) Shadow. Generated images might not include shadows, or have inconsistent shadows. iii) Objects without Support. In generated images, an object or material appears to be floating in mid-air without any visible means of support, it gives the impression that the object is defying gravity or the laws of physics.
The above cues provided by Borji are based on the observations of the failure cases of DM-generated images, while Farid analyzes the perspective [178] and lighting  [179] consistency in images synthesized by DALL·E 2. Noting that DALL·E 2 sometimes fails to maintain geometric consistency, such as parallel lines cannot converge at a common vanishing point. Additionally, while DALL·E 2 generally creates realistic lighting, there’s a tendency for the lighting direction to be more frontal or rearward relative to the camera compared with natural photographs.
As for physiological-based forensics, Borji [177] shows various physiological cues relative to eyes, teeth, ears, hair, skin, limbs, and fingers, etc, as shown in Fig. 6 a) bottom. These artifacts suggest that generative models often fall short when accurately depicting the intricate details of human extremities.
$\clubsuit$ Diffuser Fingerprints-based Methods. Every generation mechanism leaves its own unique fingerprints, which can be explored for detection, as shown in Fig. 6 b). In [95], the authors elucidated that CNNs inherently manifest certain image artifacts. These artifacts are harnessed in their Deep Image Fingerprint (DIF) methodology to extract distinct fingerprints from images generated by these networks (GANs and DMs). This approach facilitates the identification of images originating from a specific model or its fine-tuned variants. Complementarily, the studies by Wang et al. [84] and Ma et al. [176] delve into the realm of diffusion models. They have laid the groundwork for detecting DM-generated images by leveraging the deterministic reverse and denoising processes inherent diffusion models.
Wang et al. [84] propose a novel image representation called DIffusion Reconstruction Error (DIRE) based on their hypothesis that images produced by diffusion processes can be reconstructed more accurately by a pre-trained diffusion model compared to real images. DIRE measures the error between an input image and its reconstruction by a pre-trained diffusion model. The computation process of DIRE can be simply concluded as follows: the input image $\mathbf{x}_{0}$ is first gradually inverted into a noise image $\mathbf{x}_{T}$ by DDIM inversion [225] and then is denoised step by step until getting a reconstruction $\mathbf{x}_{0}^{\prime}$. DIRE is the residual image obtained from $\mathbf{x}_{0}$ and $\mathbf{x}_{0}^{\prime}$, which can be used to differentiate real or generated images.
Though Wang et al. [84] indeed leverages some deterministic aspects, their approach (i.e., DIRE) primarily targets the reconstruction at the initial time step $\mathbf{x}_{0}$. This method potentially overlooks the rich information present in the intermediate stages of the diffusion and reverse diffusion processes. Ma et al. [176] exploits these intermediate steps later. They design Stepwise Error for Diffusion-generated Image Detection (SeDID), particularly focusing on the errors between reverse and denoise samples at specific time steps during the generation process.
$\clubsuit$ Spatial-based Methods. This research collection focuses on mining spatial characteristics and features within images to detect DM-generated content. Each study utilizes different spatial aspects of images, such as texture, gradients, and local intrinsic dimensionality, for detection, see Fig. 6 c). Motivated by the principle that pixels in rich texture regions exhibit more significant fluctuations than those in poor texture regions. Consequently, synthesizing realistic rich texture regions proves to be more challenging for existing generative models. Zhang et al. [94] leverage such texture features and exploit the contrast in inter-pixel correlation between rich and poor texture regions within an image for DM-generated image forensics. Nguyen et al. [174] use gradient-based features to differentiate DM-generated and human-made artwork by Lorenz et al. [175] propose using the lightweight multi Local Intrinsic Dimensionality (multiLID) for effective detection. This approach stems from the principle that neural networks heavily rely on low-dimensional textures [226] and natural images can be represented as mixtures of textures residing on a low-dimensional manifold [227]. Leveraging this concept, multiLID scores are calculated on the lower-dimensional feature-map representations derived from ResNet18. Then a classifier (random forest) is trained on these multiLID scores to distinguish between synthetic and real images. Their proposed multiLID approach exhibits superiority in diffusion detection. However, this solution does not offer good transferability.
$\clubsuit$ Frequency-based Methods. Frequency-based methods analyze the frequency components of an image to extract information that is not readily apparent in the spatial domain, see Fig. 6 d). The initial study by Wolter et al. [171] highlights the limitations of traditional detection methods that primarily utilize spatial domain convolutional neural networks or Fourier transform techniques. In response, they propose an innovative approach using multi-scale wavelet representation, which incorporates elements from both spatial and frequency domains. However, the study notes limited benefits from higher-order wavelets in guided diffusion data, pinpointing this as a potential area for future research. In a recent study, Xi et al. [173] develop a dual-stream network that combines texture information and low-frequency analysis. This approach identifies artificial images by focusing on low-frequency forgeries and detailed texture information. It has shown to be efficient across various image resolutions and databases. [172] introduces AUSOME to detect DALL-E 2 images by performing a spectral comparison of Discrete Fourier Transform (DFT) and Discrete Cosine Transform (DCT). Bammey [70] uses the cross-difference filter on the image to highlight the frequency artifacts from DM-generated images. It shows some generalization ability, as well as robustness to JPEG compression.
3.2.2 Beyond Detection
Research on detecting DM-generated images aims not only to enhance accuracy but also to add additional functionalities to detectors. It seeks to develop a more nuanced understanding and application of these detectors, unlocking new avenues and potential in the realm of DM-generated image analysis and detection. We summarize the existing work into the following categories:
Figure 7: Illustrations of beyond detection methodologies for LAIM-generated image.
$\clubsuit$ Attribution and Model Parsing. Attribution is to recognize the specific diffusion model that generates images. Model parsing refers to the process of analyzing and interpreting the structure and characteristics of generative models. The purpose of model parsing is to gain insight into how a particular model generates images, which can be crucial for tasks like identifying the source of a synthetic image, see Fig. 7 a).
Attribution. Guarnera et al. [170] focus on attributing DM-generated images using a multi-level hierarchical approach. Each level solves a specific task: level 1 classifies images as real or AI-generated; level 2 defines whether the input images are created by GAN or DM technologies; level 3 solves the model attribution task.
Model Parsing. The model parsing task is first defined by Asnani et al. [91] as estimating Generative Model (GM) network architectures (e.g., convolutional layers, fully connected layers, and the number of layers) and their loss function types (e.g., Cross-entropy, Kullback–Leibler divergence) by examining their generated images. To tackle this problem, they compiled a dataset comprising images from 116 GMs, which includes detailed information about the network architecture types, structures, and loss functions utilized in each GM. They introduce a framework with two components: a Fingerprint Estimation Network, which estimates a GM fingerprint from a generated image, and a Parsing Network, which predicts network architecture and loss functions from the estimated fingerprints.
$\clubsuit$ Generalization. A generalizable detector can successfully detect the unseen images generated by the newly released generators, as shown in Fig. 7 b).
Epstein et al. [163] collect a dataset generated by 14 well-known diffusion models and simulate a real-world learning setting with incremental data from new DMs. They find that the classifiers generalize to unseen models, although performance drops substantially when there are major architectural changes. Different from [163], [89, 164] utilize pre-trained models, ViT [139] and CLIP [138] respectively, to achieve high generalization and robustness in various scenarios.
In addition, [165] uses data augmentation approaches to improve detection generalization. To avoid data dependency on particular generative models and improve generalizability, Jeong et al. [166] propose a framework that can generate various fingerprints and train the detector by adjusting the amplitude of these fingerprints or perturbations. Their method outperforms the prior state-of-the-art detectors, even though only real images are used for training. Reiss et al. [167] introduce FACTOR, particularly for fact-checking tasks. Although their method is training-free, it achieves promising performance. Dogoulis et al. [168] address the generalization of detectors in cross-concept scenario (e.g., when training a detector on human faces and testing on synthetic animal images), and propose a sampling strategy that considers image quality scoring for sampling training data. It shows better performance than existing methods that randomly sample training data. DNF [169] seeks to address generalization challenges through an ensemble representation that estimates the noise generated during the inverse diffusion process and achieves state-of-the-art effects of detection.
$\clubsuit$ Interpretability.This picture looks like someone I know, and if the AI algorithm tells it is fake or real, then what is the reasoning and should I trust? " Detection with interpretability is working towards solving such question, see Fig. 7 c). In pursuit of creating interpretable detectors, Aghasanli et al. [162] propose a method for DM-generated image detection, taking advantage of features from fine-tuned ViTs combined with existing classifiers such as Support Vector Machines.
$\clubsuit$ Localization. A detector that can localize the artifacts in a DM-generated image or give a prediction with a localization map indicating which input regions have been manipulated, as shown in Fig. 7 d).
Fully-supervised. Localization in a fully-supervised setting requires detectors to be explicitly trained for localization as a segmentation problem with localization mask label. Guo et al. [83] propose a hierarchical fine-grained Image Forgery Detection and Localization (IFDL) framework with three components: a multi-branch feature extractor, localization, and classification modules. The localization module segments pixel-level forgery regions, while the classification module detects image-level forgery. [161] centers on detecting and segmenting artifact areas that are only noticeable to human perception, not full manipulation region. In addition to providing a pixel-level localization map, the study by Guillaro et al. [87] also offers an integrity score to assist users in their analysis.
Weakly-supervised. In contrast to the above work, which addresses localization in a fully-supervised setting. Tantaru et al. [160] consider a weakly-supervised scenario motivated by the fact that ground truth manipulation masks are usually unavailable, especially for newly developed local manipulation methods. They conclude that localization of manipulations for latent diffusion models [81] is very challenging in the weakly-supervised scenario.
$\clubsuit$ Robustness. A robust detector is strategically developed to counteract different attacks consisting of intentionally designed perturbations or noise, as shown in Fig. 7 e). It is also designed to maintain consistent detection performance when exposed to real-world conditions, such as dealing with image compression and blurring.
Spatial-based. Ju et al. [157] propose a synthesized image dataset that can contain diverse post-processing techniques to simulate real-world applications. They further introduce the Global and Local Feature Fusion Framework to enhance image detection by integrating multi-scale global features with detailed local features. Another work presented by Xu et al. [158] also leverages multi-level feature representation for enhancing robustness. Different from the above works, [159] highlights the effectiveness of computing local statistics, as opposed to global statistics, in distinguishing digital camera images from DM-generated images. This method produces promising robustness to image resizing and JPEG compression.
Frequency-based. Contrasting with the approaches that mine spatial features, Hooda et al. [156] leverage frequency features through a unique disjoint ensembling approach. They utilize redundant information in the frequency feature space of DM-generated images to generate disjoint ensembles for adversarial DM-generated image detection.
$\clubsuit$ Empirical Study. The empirical study serves as a crucial foundation for devising methods to detect images generated by rapidly advancing DMs. Despite DMs’ rapid progression, detection methods have not evolved at the same pace. Therefore, thorough experiments and insights from these empirical studies are vital for advancing detection technologies, as illustrated in Fig. 7 f).
Recent studies by Corvi et al. [93] and Ricker et al. [88] investigate the challenge of differentiating photorealistic synthetic images from real ones. They find that detectors trained only on GAN images work poorly on DM-generated images. Additionally, DMs exhibit much subtler fingerprints in the frequency domain than GANs.
Realizing it is challenging for existing methods to detect DM-generated images, Corvi et al. [153] extend their previous work [93] to gain insight into which image features better discriminate fake images from real ones. Their experimental results shed light on some intriguing properties of synthetic images:
  • To date, no generator appears to be artifact-free. Unnatural regular patterns are still observed in the autocorrelation of noise residuals;
  • When a model is trained on a limited-variety dataset, its biases can transfer to the generated images;
  • Synthetic and real images differ notably in their mid-high frequency signal content.
These traces exploited in this empirical study can be instrumental in developing forensic detectors.
Cocchi et al. [110] investigate the robustness of different pre-trained model backbones against image augmentations and transformations. Their findings reveal that these transformations significantly affect classifier performance.
Other works [154, 62] empirically study human face detection. [154] focuses on detecting profile photos of online accounts, finding that a detector (EfficientNet-B1 [228]) only trained on face images completely fails when detecting non-face images, and the detector can learn the semantic-level artifacts because it remains highly accurate even when the image is compressed to a low resolution. [62] investigates the difficulty of non-expert humans in recognizing fake faces and further concludes that trained models outperform non-expert human users, which brings out the need for solutions to contrast the spread of disinformation.
Furthermore, [67] centers on the biological images and narrows on detecting synthetic western blot images. They explore a range of detection strategies, such as binary classification and one-class detectors, and notably, they do this without utilizing synthetic images in training. Their findings demonstrate the successful identification of synthetic biological images. [155] is the first work to detect authentic-looking handwriting generated by DMs. Their experiments reveal that the strongest discriminative features come from real-world artifacts in genuine handwriting that are not reproduced by generative methods.
3.3 Video
Recently, there has been limited research focused on detecting videos generated by LAIMs. This is primarily because video generation techniques are more complex than those for images. Additionally, as outlined in Section 2, video generation technology is still in its nascent stages compared with image generation. To the best of our knowledge, only one work [152] related to this area, see Fig. 8, and we classify it into the Beyond Detection category.
Figure 8: Illustration of detection methodology in generalization task for LAIM-generated video.
3.3.1 Beyond Detection
$\clubsuit$ Generalization. Many works address generalization in deepfake video detection. However, they may overlook one challenge: the standard out-of-domain evaluation datasets are very similar in form to the training data, failing to keep up with the advancements in DM video generation. [152] addresses such an issue by introducing the Simulated Generalizability Evaluation (SGE) method, which involves simulating spatial and temporal deepfake artifacts in videos of human faces using a Markov process. SGE aims to improve the generalizability of video detection by creating artifacts that reflect implausibilities in facial structures, which could accompany unseen manipulation types.
3.4 Audio
The research on the detection of LAIM-generated audio is also very limited. The reason is that visual and text content are more dominant and widespread in media and online platforms. This prevalence makes fake text/images more common and thus a higher priority for detection. However, as voice synthesis and manipulation technologies improve and become more accessible, the importance and focus on detecting fake audio are likely to increase.
Figure 9: The artifacts introduced by DM-based neural vocoders (WaveGrad [134] and DiffWave [105]) to a voice signal [103]. The differences in mel-spectrograms between real and generated ones are illustrated in the third and fifth columns.
3.4.1 Pure Detection
$\clubsuit$ Vocoder-based. Sun et al. [103] focus on detecting unique artifacts generated by neural vocoders in audio signals. Visible artifacts introduced by different neural vocoder models can be observed in Fig. 9. The study employs a multi-task learning framework incorporating a binary-class RawNet2 [229] model. This model uniquely shares a feature extractor with a vocoder identification module. By employing this shared structure, the feature extractor is effectively guided to concentrate on vocoder-specific artifacts. However, in their method, the detector may become overly specialized in detecting specific forgery technologies, hindering generalization to unseen vocoder artifacts.
3.5 Multimodal
Here, we categorize literature utilizing multimodal learning, which takes multiple modalities for detecting forgeries in single or multiple modalities. In particular, multimodal learning [137] refers to an embodied learning situation that includes learning multiple data modalities such as text, image, video, and audio.
Figure 10: Illustrations of pure detection methodologies for LAIM-generated multimodal media.
3.5.1 Pure Detection
$\clubsuit$ Text-assisted. This type of method utilizes text-image or prompt-image pair as input and focuses on extracting and learning features from both text and images to improve DM-generated image detection, see Fig. 10 a). Amoroso et al. [151] leverage the semantic content of textual descriptions alongside visual data. They introduce a contrastive-based disentangling strategy to analyze the role of the semantics of textual descriptions and low-level perceptual cues. The experiments are conducted on their proposed COCOFake dataset, where CLIP [138] is employed as the backbone for feature extraction. These extracted features are then used to train a logistic regression model. They find that the visual features extracted from the generated image still retain the semantic information of the original caption used to create it, which allows them to distinguish between natural images and the generated ones using only semantic cues while neglecting the perceptual ones.
$\clubsuit$ Text-image Inconsistency. This refers to a type of misinformation where the textual content does not align with or accurately represent the original meaning or intent of the associated image. This type of misinformation is commonly presented in the guise of news [230] to mislead the public, see Fig. 10 b). Tan et al. [13] create a NeuralNews dataset which contains articles with LLM-manipulated text and authentic images tailored for such task. They further propose DIDAN (Detecting Cross-Modal Inconsistency to Defend Against Neural Fake News) framework that effectively leverages visual-semantic inconsistencies between article text, images, and captions, offering a novel approach to counteract neural fake news.
Figure 11: Illustrations of beyond detection methodologies for LAIM-generated multimodal media.
3.5.2 Beyond Detection
$\clubsuit$ Attribution. This refers to research works that utilize a multimodal learning approach, combining prompt information with image features, to enhance the accuracy of attributing DM-generated images to their source model, shown in Fig. 11 a). Sha et al. [150] build an image-only detector (ResNet18) and a hybrid detector (take advantage of CLIP’s [138] image and text encoders for feature extraction). Empirical results show that learning images with prompts achieves better performance than image-only attribution regardless of the dataset.
$\clubsuit$ Generalization. This area of research focuses on developing a fake image detector that is guided by language, enhancing its ability to detect a wider range of new, unseen DM-generated images, shown in Fig. 11 b).
Prompt Tuning. In this vein, Chang et al. [149] employ soft prompt tuning on a language-image model to treat image detection as a visual question-answering problem. These prompt questions (e.g., “Is this photo real?”) are then converted into vector representations, later are fed into Q-former and LLM with the image features extracted by the image encoder, therefore the DM-generated image detection problem is formulated as a visual question-answering problem, the LLM gives the detection output “Yes” for real images and “No” for fake images.
Contrastive Learning. [148] adopt a language-guided contrastive learning approach, augmenting training images with textual labels (e.g., “Real/Synthetic Photo” and “Real/Synthetic Painting”) for forensic feature extraction.
$\clubsuit$ Interpretability. It provides evidence with detection results by leveraging multimodal media, shown in Fig. 11 c). To combat misinformation and emphasize the significance of scalability and explainability in misinformation detection, Xu et al. [147] propose a conceptual multi-modal framework composed of 4 levels: signal, perceptual, semantic, and human, for explainable multimodal misinformation detection.
$\clubsuit$ Localization. Detection methods with localization aim to not only detect the authenticity of media content but also locate the manipulated content (i.e., image bounding boxes and text tokens). Unlike leveraging multimodal learning for enhanced single-modal forgery detection tasks, the methods we summarize here can detect and locate multi-modal media, shown in Fig. 11 d).
Spatial-based. Shao et al, [14] initiate a significant development in the field of multimodal media manipulation detection with their dataset “Detecting and Grounding MultiModal Media Manipulation" ($DGM^{4}$), where image-text pairs are manipulated by various approaches, to cope with the threat that enormous text fake news is generated or manipulated by LLMs. They further use contrastive learning to align the semantic content across different modalities (image and text) for detection and localization. Subsequent research [231, 145, 146] in this area has built upon this dataset. Specifically, [231] is an extension of [14] by integrating Manipulation-Aware Contrastive Loss with Local View and constructing a more advanced model HAMMER++ [231], which further improves detection and localization performance. Wang et al. [146] construct a simple and novel transformer-based framework for the same task. Their designed dual-branch cross-attention and decoupled fine-grained classifiers can effectively model cross-modal correlations and exploit modality-specific features, demonstrating superior performance compared to HAMMER [14] and UFAFormer[145] on $DGM^{4}$.
Frequency-based. UFAFormer (a Unified Frequency-Assisted transFormer framework) [145] incorporates frequency domain to address the detection and localization problem. It simultaneously integrates and processes the detection and localization processes for manipulated faces, text, as well as image-text pair, simplifying the architecture in [14] and facilitating the optimization process.
$\clubsuit$ Empirical Study It provides insights into the feasibility of detecting generated images by leveraging multimodal media shown in Fig. 11 e). [143] shows that it is possible to detect the generated images using simple Multi-Layer Perceptrons (MLPs), starting from features extracted by CLIP [138] or traditional CNNs. They found that incorporating the associated textual information with the images rarely leads to significant improvement in detection results but that the type of subject depicted in the image can significantly impact performance. Papadopoulos et al. [144] highlight that specific patterns and biases in multimodal misinformation detection benchmarks can result in biased or unimodal models outperforming their multimodal counterparts on a multimodal task. Therefore they create a ”VERification of Image-TExt pairs” (VERITE) benchmark to address the unimodal bias in multimodal misinformation detection.
4 Tools
In this section, we present a selection of popular, user-friendly detection tools we have evaluated, see Table II.
$\clubsuit$ Text. In the domain of LLM-generated text detection tools, diverse instances illuminate specific strengths and drawbacks. Copyleaks [232] offers high accuracy across 30 languages but faces performance delays and lacks text exclusion options. ZeroGPT [233] is user-friendly but doesn’t support file analysis. Though Winston AI[234] presents a free OCR (Optical Character Recognition)-powered alternative with an accessible interface, it needs enhancements in reporting and language support. Simplicity and speed take center stage with Crossplag’s[235], but it is confined to English and lacks graphical elements. GLTR[236] delivers in-depth, color-coded text analysis but is less intuitive for beginners and struggles with newer technologies. Content at Scale[237] yields easy-to-interpret results but may exhibit inconsistencies in AI rewrites. Undetectable.ai [238] supports various models but may discourage users due to pricing concerns. Illuminarty[239] excels in detecting AI-generated texts but introduces complexity and slower processing for basic needs. Finally, Is it AI[240] matches other tools in function but faces reliability issues, especially with short or non-English texts.
Modality Tool Company Link Reference
AI Content Detector Copyleaks Link [232]
AI Content Detector, ChatGPT detector ZeroGPT Link [233]
AI Content Detector Winston AI Link [234]
AI Content Detector Crossplag Link [235]
Giant Language model Test Room GLTR Link [236]
The AI Detector Content at Scale Link [237]
AI Checker Originality ai Link [241]
Advanced AI Detector and Humanizer Undetectable ai Link [238]
AI Content Detector Writer Link [242]
AI Content Detector Conch Link [243]
Illuminarty Text Illuminarty Link [239]
Text AI-Generated Text Detector Is it AI Link [240]
AI or Not image AI or Not Link [244]
AI-Generated Image Detector Is it AI Link [245]
Illuminarty Image Illuminarty Link [246]
SynthID Google Link [247]
Advanced AI Image Detector Content at Scale Link [248]
Image AI Image Detector Huggingface Link [249]
Audio AI or Not audio AI or Not Link [244]
TABLE II: Existing popular detection tools for exposing LAIM-generated multimedia.
$\clubsuit$ Image. In evaluating tools for detecting images generated by DMs, AI or Not [244] demonstrated fast execution while identifying fake images and supported diverse formats. However, it cannot provide stable results. In addition, Is it AI[245] and Content at Scale[248] offer a user-friendly interface with limitations in detecting certain content types. Illuminarty[246] provides in-depth analysis but faces usability complexities, and Huggingface[249] excelled in accessibility with limitations in advanced analysis. Google has developed SynthID[247], specifically designed for images generated by Imagen[250].
$\clubsuit$ Audio. AI or Not[244] also provides an AI audio detection tool which sets it apart from other AI detection tool product lineups. Currently, they support WAV, MP3, and FLAC audio files of a maximum size of 100MB and have a duration of more than 5 seconds.
5 Discussion
In this section, we discuss the main challenges and limitations encountered in the era of detection based on our comprehensive review of existing works. Then we propose potential directions for future research aimed at developing more efficient and practically effective detection systems.
5.1 Challenges
$\clubsuit$ Limited Reliability of Detection Models. A reliable detector should retain promising generalizability to new emerging models or data sources of detection, good interpretability in building trust in prediction results, and robust resistance against different real-life application conditions [251]. Therefore, we pinpoint the reliability challenges in detection models from three aspects: generalizability, interpretability, and robustness.
  • Generalizability. We can see from Fig. 1, that LAIMs and their generated multimedia are getting actively updated every three months. However, many current detectors are tailored to learn specific features from their training data, leading to overfitting and a lack of flexibility in detecting unseen data. For example, [44, 182, 183, 48, 184, 163] all highlight such generalization issues in detectors, especially in  [44, 163], both of them aim to simulate real-world scenarios. [44] demonstrates that commonly-used detectors have difficulty in detecting text in the wild. Similarly,  [163] shows that the detector’s performance drops substantially when the testing images come from generators that have major architectural changes from those that generate training data. All of those findings reveal the difficulty of reliable generalization to new emerging models or data sources of detection.
  • Interpretability. As multimedia generated by LAIMs becomes increasingly realistic and indistinguishable from human-created content, there is a growing reliance on methods that are easily understandable and tractable [252] instead of those with high accuracy but derived based on data that are hard to explain. Efforts like [190, 38] employ SHAP [221] to analyze local explanations of model’s decisions. However, their explanations are data-dependent.  [162] employs prototype-based explanation method, which can be susceptible to spurious correlations (confounding bias) [253]. The above methods still do not accurately reflect a detector’s decision-making process. Besides the model architecture, training data is another challenging aspect to consider, particularly in determining which specific features or subsets of the vast training data contribute to a detector’s predictions.
  • Robustness. Potential attacks are an important factor in the continued unreliability of current LAIM-generated multimedia detectors. These detectors are tasked with not only identifying content produced by the latest LAIMs but also demonstrating resilience against sophisticated attacks created by the most advanced generative models. [254, 255] illustrate this vulnerability, where they use diffusion-based models to craft attacks that successfully exploit weaknesses in existing detectors. Furthermore, [110] has shown that the performance of detectors is significantly influenced by image transformations. This finding emphasizes the challenges in ensuring the robustness of detectors, especially in real-world scenarios where images often undergo compression or blurring. Additionally, LAIM-generated text detection faces paraphrase attacks and prompt attacks performed by LAIMs, besides adversarial attacks [18, 19]. In a nutshell, LAIM-generated attacks and real-life detection scenarios bring a greater challenge to the robustness of detectors ever than before.
$\clubsuit$ Constrained Size, Accessibility, Range, and Variety of Datasets. As illustrated in Table I, there is a significant gap in the availability of large-scale datasets for video, audio, and multimodal detection. While there are a few million-scale image detection datasets, many have limited sample sizes (some are even less than ten thousand). Additionally, there is a noticeable lack of diversity in datasets. Current public multimodal datasets focus primarily on image and text modalities, overlooking valuable information that could be derived from other modalities like video and audio. Furthermore, the majority of text detection datasets are predominantly in English, and are limited to specific domains without encompassing a broad range of topics. The available fake audio dataset, LibriSeVoc [103], contains only two diffusion-based vocoders. This lack of diversity and narrow range lead to significant out-of-distribution challenges, causing detectors to underperform markedly when faced with data in real-world settings.
$\clubsuit$ Lack Evaluation Benchmark. The field of LAIM-generated multimedia detection faces the challenge of lacking a standardized benchmark, leading to inconsistent data processing, evaluation strategies, and metrics. This results in unfair comparisons and potentially misleading outcomes. For instance, Kamat et al. [152] note that detectors often tested on datasets similar to their training data, while Cocchi et al. [110] demonstrates that different data augmentation has notable impacts on the performance of detectors. Furthermore, the capability of LAIMs to partially manipulate multimedia content, such as through image inpainting, polishing human-written text, and synthesizing specific video frames, adds to the evaluation complexity. Therefore, how best to evaluate such intricately manipulated content with a unified benchmark remains challenging.
5.2 Future Directions
$\clubsuit$ Building Foundation Models. The term “foundation models” was first introduced by Bommasani et al. [256], and they defined foundation model as “the base models trained on broad data that can be adapted to a wide range of downstream tasks”, (e.g., CLIP [138], DALL-E 3 [120], ViT [139], and GPT family) . Multimedia generation has recently witnessed remarkable progress fueled by foundation models [4]. In LAIM-generated text detection, foundation models are widely used. Most recently, in fake image detection [150, 151, 149, 143, 162, 164], the paradigm is also shifting towards foundation models, because it allows rapid model development and provides better performance in real-world scenarios with their extensive pre-training on diverse datasets.
However, the application of foundation models in detection tasks remains quite limited. For instance, the above efforts in synthesized image detection are primarily limited to CLIP [138] and ViT [139]. This limitation is even more pronounced in areas like audio, video, and multimodal forgery detection, where there is a notable scarcity of research utilizing or developing foundation models. This gap indicates a significant opportunity for advancement in the field, particularly in integrating and exploring existing foundation models to bolster the detection capabilities of a wide range of digital forgeries. More importantly, there is a pressing need to develop multimodal detection-oriented foundation models. Such models would offer more robust and versatile detection systems compared to task-specific models, enabling researchers to rapidly adapt them for various downstream detection tasks, streamlining the detection process, and enhancing its efficacy.
$\clubsuit$ Harnessing Generative Discrepancies. Recent studies reveal clear differences between LAIM-generated multimedia and human-created content. LAIM-generated text, as shown by [50, 257, 258], often exhibits longer, more deterministic, and complex syntax compared to human writing. In image generation, current DMs face challenges in producing high-frequency information[88, 171], rich textures [94], repetitive patterns [177], and adherence to physical principles like lighting [179]. Effective detection strategies might involve a mix of these indicators, with potential in ensemble learning  [259], foundation models  [256], multi-modal/task learning, and cross-modal feature embedding.
$\clubsuit$ Ensuring Reliability of Detectors. As reliability challenges in detection methods are discussed above, we urge forthcoming detection studies to prioritize detector reliability. The development of a reliable detector should aim for a unified framework that ensures generalizability, robustness, and interpretability concurrently. This could be a promising direction for detectors to combat LAIM-generated multimedia in real-world scenarios and increase the trust of the social community in detection results. Training on a diverse and large-scale dataset, building robust detectors with attacks, incorporating explainable AI (XAI) techniques [252], considering strategies like online learning [163] or transfer learning with new data to keep the model updated would contribute to building a reliable detector.
$\clubsuit$ Establishing Fair Detectors. It has been revealed that current detectors, though with high detection accuracy, result in unfair performance disparities among demographic groups, such as race and gender [260, 261]. This can lead to particular groups facing unfair targeting or exclusion from detection, potentially allowing misclassified LAIM-generated multimedia to manipulate public opinion and undermine trust in the model. Despite its criticality, the issue of fairness in LAIM-generated multimedia detection, even in detecting deepfake generated by traditional methods, has not received adequate attention from the research community. We advocate for a proliferation of studies that delve into addressing the bias problem in detectors. This involves not only developing a fair detector, but also ensuring that this fairness is generalizable across various scenarios without showing bias against certain groups.
$\clubsuit$ Accelerating the Development of Multimodal Detectors and Datasets. Existing research on detecting LAIM-generated multimedia mainly focuses on single data modality. However, in real-world scenarios, it is very hard to know which data modality is generated. For example, generated videos may have both visual and audio components generated, or only one modality may be created. Hence, the creation of multimodal detectors is crucial. To effectively develop these models, assembling extensive multimodal datasets for their training is a fundamental requirement. This can be explored in the future.
$\clubsuit$ Towards Interpretable User-friendly Detection Tools. Our analysis of popular online detection tools discussed in Section 4, reveals a common shortcoming: a lack of interpretability. However, humans generally prefer tools that are directly interpretable, tractable, and trustworthy [252]. Developing interpretable detection tools can gain more social attention and popularity. Therefore, enhancing interpretability is not just a technical upgrade but a significant step in making these tools more accessible to broader users, especially beneficial for underrepresented groups such as non-technical persons and elderly adults.
6 Conclusion
This paper provides the first systematic and comprehensive survey covering existing research on detecting multimedia from text, image, audio, and video to multimodal content generated by large AI models. We introduce a novel taxonomy for detection methods within each modality, categorizing them under two primary frameworks: pure detection (focusing on improving detection accuracy) and beyond detection (integrating attributes like generalizability, robustness, and interoperability to detectors). Additionally, we have outlined the sources contributing to detection, such as generation mechanisms of LAIMs, public datasets, and online tools. Finally, we pinpoint current challenges in this field and propose potential directions for future research. We believe that this survey serves as the initial contribution to addressing a notable academic gap in this field, aligning with global AI security initiatives, thereby upholding the authenticity and integrity of digital information.